OpenStudy (anonymous):

A bacteria culture is grown with a population of 45 million cells and kept under observation. The cell population is found to double every hour. What will the population be after four hours? 96 million 124 million 360 million 720 million

OpenStudy (anonymous):

help please @iGreen

OpenStudy (igreen):

The population doubles every hour. 45 * 2 = 90 90 * 2 = 180 180 * 2 = 360 360 * 2 = ?

OpenStudy (igreen):

@brittany.hodges What's 360 * 2? That's your answer.

OpenStudy (anonymous):

thanks gosh I love you right know you have no idea how big of a help that was

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

?

OpenStudy (igreen):

You're welcome! :)